### v2

@@ 1 Introduction
Let the analyst’s model be an incomplete information game with two players $1$ and $2$; two actions $\alpha$ and $\beta$ for each player; and two payoff-relevant states $\theta_{1}$ and $\theta_{2}$ occurring with equal probability of 1/2121/21 / 2.111This game is discussed by Liu (2015). Each player’s set of types is a singleton; that is, it is common knowledge that each player has a uniform prior over $\{\theta_{1},\theta_{2}\}$. The payoffs are summarized in Table 1, where players would be better coordinated by choosing the same actions in state $\theta_{1}$ and the different actions in state $\theta_{2}$. This game has an infinite number of Bayes Nash equilibria (BNE) because the expected payoff for each player is $1/2$ under every action profile.
@@
$\alpha$ 00 $1/4$
$\beta$ $1/4$ 00
Table 2: Joint probabilities of actions and a state
@@
A generalized potential function is associated with a belief-invariant correlating device that assigns each player an action subset as a recommendation to choose an action from that subset. An important belief-invariant correlating device is one designed to maximize the expected value of a generalized potential function. For any such correlating device, there exists a BIBCE in which each player chooses an action from the recommended subset. This BIBCE is referred to as a GP-maximizing BIBCE.
@@
The main result is applied to supermodular games in Section 6. We address the question of when a BNE can be robust and show that if a game is a supermodular pontential game with a unique potential-maximizing BNE, then this BNE is robust. As a numerical example, we consider a binary-action supermodular game with two players, which is a global game (Carlsson and van Damme, 1993) with discretized signals and no dominance regions. This game has a robust BNE, where players always choose ex-post risk-dominant actions. Even in the absence of dominance regions, allowing a richer set of perturbations for the incomplete information game leads to the same outcome as the global game selection.
@@
The concept of BIBCE is introduced by Forges (2006) and Lehrer et al. (2010) under the additional restriction that action recommendations cannot depend on the state. A BIBCE is a special case of a BCE, which is introduced by Bergemann and Morris (2013, 2016) as an analyst’s prediction when she cannot rule out players having additional information. The work of Bergemann and Morris (2013, 2016) also supports the use of BIBCE in our paper as a candidate for an analyst’s prediction under insufficient and imprecise information. [^2] For further motivation to study BIBCE and its relationship to BCE, see Liu (2015) and Bergemann and Morris (2017). For its relation to other special cases of BCE, see Bergemann and Morris (2019).
@@ 2.1 Incomplete information games
  • $\Theta=\prod_{i\in I}\Theta_{i}$ is an at most countable set of payoff-relevant states, where player $i$’s payoff function is determined by the $i$-th component $\theta_{i}\in\Theta_{i}$.
@@
Payoff functions are assumed to be bounded, i.e., $\sup_{i,a,\theta}|u_{i}(a,\theta)|<\infty$. For each $i\in I$, let $T_{i}^{*}\subseteq T_{i}$ and $\Theta_{i}^{*}\subseteq\Theta_{i}$ denote the sets of player $i$’s types and payoff-relevant states on the support of $\pi$, respectively: $T_{i}^{*}=\{t_{i}\in T_{i}\mid\pi(t_{i})>0\}$ and $\Theta_{i}^{*}=\{\theta_{i}\in\Theta_{i}\mid\pi(\theta_{i})\}>0$, where $\pi(t_{i})\equiv\sum_{t_{-i},\theta}\pi(t,\theta)$ and $\pi(\theta_{i})\equiv\sum_{t,\theta_{-i}}\pi(t,\theta)$ are the marginal probabilities. Player $i$’s belief is given by $\pi(t_{-i},\theta|t_{i})\equiv\pi(t,\theta)/\pi(t_{i})$ when his type is $t_{i}\in T_{i}^{*}$.
@@ 2.3 Elaborations
Note that if $(\bar{T},\Theta,\bar{\pi},u)$ is the conjunction of $(T,\Theta,\pi,u)$ and a belief-invariant communication rule $\rho$, it is an elaboration of $(T,\Theta,\pi,u)$ with $\tau_{i}$ and $\mu_{i}$ given by $\tau_{i}(t_{i},m_{i})=t_{i}$ and $\mu_{i}(t_{i},m_{i})=m_{i}$ for all $\bar{t}_{i}=(t_{i},m_{i})\in\bar{T}_{i}$ and $i\in I$ [^7] Equation (1) implicitly requires that $\sum_{\bar{t}\in\tau^{-1}(t)}\rho(\mu(\bar{t})|t,\theta)=1$ for all $(t,\theta)\in T^{*}\times\Theta^{*}$ and $\tau(\bar{T}^{*})=T^{*}$ because $\bar{\pi}(\bar{T}\times\Theta)=1$.
@@
A game $(\bar{T},\Theta,\bar{\pi},u)$ is an elaboration of $(T,\Theta,\pi,u)$ if and only if, for each $i\in I$, there exists a mapping $\tau_{i}:\bar{T}_{i}\to T_{i}^{*}$ such that
@@ 2.4 Non-redundant games
A non-redundant game [^9] We can characterize a non-redundant game using belief hierarchies. Let $[t_{i}]\equiv\{t_{i}^{\prime}\in T_{i}\mid h_{i}(t_{i})=h_{i}(t_{i}^{\prime})\}$ and $[t]=([t_{i}])_{i\in I}$. Consider $(T^{0},\Theta,\pi^{0},u)$ with $T^{0}=\{[t]\mid t\in T\}$ and $\pi^{0}([t],\theta)=\pi(\{(t^{\prime},\theta)\,|\,t^{\prime}\in[t]\})$. Then, it is straightforward to show that $(T^{0},\Theta,\pi^{0},u)$ is a non-redundant game of $(T,\Theta,\pi,u)$. will serve as the analyst’s model in the next section. Liu (2015) shows that if two games have the same set of belief hierarchies and one is non-redundant, then the other is an elaboration of the non-redundant game. [^10] Liu (2015) assumes a finite type space.
@@ 3 Robustness
We further assume that the analyst believes that the true game is in a “neighborhood” of elaborations of the analyst’s model. We call such a nearby game an $\varepsilon$-elaboration, which is approximately an elaboration of $(T,\Theta,\pi,u)$. An elaboration is a special case of $\varepsilon$-elaboration with $\varepsilon=0$.
@@
    for all $\bar{t}_{i}\in\bar{T}_{i}^{\flat}$, where $\tau(\bar{t})=(\tau_{i}(\bar{t}_{i}))_{i\in I}$ and $\tau_{-i}(\bar{t}_{-i})=(\tau_{j}(\bar{t}_{j}))_{n\neq i}$. We call the mapping $\tau$ an elaboration mapping (with some abuse of terminology).
@@
for some $\sigma\in\mathcal{E}$, where $\bar{\sigma}\circ{\bar{\pi}}(a,\bar{t},\theta)=\bar{\sigma}(a|\bar{t},\theta)\bar{\pi}(\bar{t},\theta)$ and $\sigma\circ\pi(a,t,\theta)=\sigma(a|t,\theta)\pi(t,\theta)$. If =$\mathcal{E}=\{\sigma\}$ is a singleton, then $\sigma$ is said to be robust.
To interpret (7), recall Lemma 4, which says every 00-elaboration has a BIBCE $\bar{\sigma}$ that is equivalent to a BIBCE $\sigma$ of $(T,\Theta,\pi,u)$, where (4) holds for all $(a,t,\theta)\in A\times T\times\Theta$. Equation (7) means that (4) holds approximately for sufficiently small $\varepsilon>0$ in the case of $\varepsilon$-elaborations.
@@
Given the definition of robustness, the motivating example in Section 1 is summarized as follows and illustrates why we focus on BIBCE in our robustness concept. We have shown that there exists an $\varepsilon$-elaboration with a unique BNE. This BNE is close to the BIBCE in Table 2 when $\varepsilon$ is small, but it is not close to any BNE of the analyst’s model. This implies that the set of BNE for the analyst’s model is not robust. This is why we adopt BIBCE as the equilibrium concept for the analyst’s model. We also adopt BIBCE as the equilibrium concept for $\varepsilon$-elaborations. This is because there exists an $\varepsilon$-elaboration whose BNE is not the BIBCE in Table 2. For instance, the analyst’s model is a 00-elaboration, where every BNE differs from the BIBCE in Table 2.
@@
In Kajii and Morris (1997), the analyst’s model is a complete information game, and the solution concept for the model is a correlated equilibrium. A correlated equilibrium is defined to be robust if, for sufficiently small $\varepsilon>0$, every $\varepsilon$-elaboration has a BNE that is close to the correlated equilibrium. See Section 7.1 for further details.
@@ 4.2 Generalized potentials and BIBCE
A generalized potential function (Morris and Ui, 2005) is defined as a function over $\mathcal{A}\times\Theta$, which contains information about players’ preferences.
@@
for all $i\in I$, $a\in A$, and $\theta\in\Theta^{*}$. The next lemma shows that a potential function is a special case of a generalied potential function.
@@ 5.1 Robustness of GP-maximizing BIBCE
Every game admits a generalized potential function with the domain $\mathcal{A}=\{\{A\}\}$, in which case $\mathcal{E}^{F}$ is the set of all BIBCE. Thus, if a BIBCE is unique, which is the only element of $\mathcal{E}^{F}$, it is robust.
@@ 5.2 Proof of Theorem 1
Now, let ¯$\overline{\mathcal{E}}^{F}$ be the set of all GP-maximizing BIBCE of $(\bar{T},\bar{\Theta},\bar{\pi},\bar{u})$. Then, by Lemma 8, $\{(\bar{T},\bar{\Theta},\bar{\lambda}^{k},\bar{u})\}_{k=1}^{\infty}$ has a sequence of BIBCE $\{\bar{\sigma}^{k}\}_{k=1}^{\infty}$ such that
@@ 6.1 Supermodular potential games
Let $A_{i}$ be linearly ordered with $\geq_{i}$. We write $\geq_{I}$ for the product order: for $a,b\in A$, $a\geq_{I}b$ if and only $a_{i}\geq_{i}b_{i}$ for all $i\in I$. We say that $(T,\Theta,\pi,u)$ is a supermodular game if, for each $\theta\in\Theta$, the ex-post game with a payoff function profile $(u_{i}(\cdot,\theta))_{i\in I}$ exhibits strategic complementarities; that is, for each $i\in I$ and $a,b\in A$ with $a\geq_{I}b$,
@@
For each $\sigma\in\Sigma^{BI}$, we define a pure strategy profile $\overline{\sigma}=(\overline{\sigma}_{i})_{i\in I}$, where for each $i\in I$ and $t_{i}\in T_{i}$, player $i$ of type $t_{i}$ chooses the maximum action $a_{i}$ such that $\sigma_{i}(a_{i}|t_{i})=\sigma\left(\{a_{i}\}\times A_{-i}|\left(t_{i},t_{-i}\right),\theta\right)>0$. Similarly, we define another pure strategy profile $\underline{\sigma}=(\underline{\sigma}_{i})_{i\in I}$, where player $i$ of type $t_{i}$ chooses the minimum action that satisfies the same condition.
The following proposition establisehs that if $\sigma$ is a P-maximizing BIBCE, then the strategy profiles $\overline{\sigma}$ and $\underline{\sigma}$ are also P-maximizing BIBCE.
@@
As an example of such a BNE, we consider a binary-action supermodular game with two players, each choosing either action 1 or action 0 (which can be interpreted as “invest” and “not invest,” respectively). The set of states is given by $\Theta=\{0,1,2,3,\ldots\}$. A payoff table and a potential function for each state $n\in\Theta$ are given below, where $10<r<$ 1.
@@
action 1 ${r}^{n+1}$ 00
action 0 00 $1-{r}^{n+1}$
State $n\in\Theta$ occurs with probability $p(1-p)^{n}>0$, where $p\in(0,1)$. Each player has the following information partition of $\Theta$: player $1$’s partition is $\{\{0\},\{1,2\},\{3,4\},\ldots\}$, and player $2$’s partition is $\{\{0,1\},\{2,3\},\{4,5\},\ldots\}$, which are analogous to those in the email game. The type of a player with a partition $\{k-1,k\}$ (or $\{0\}$) is referred to as type $k$ (or type 00). Thus, the sets of player 1’s types and player 2’s types are $T_{1}=\{0,2,4,\ldots\}$ and $T_{2}=\{1,3,5,\ldots\}$, respectively. This game can be interpreted as a global game (Carlsson and van Damme, 1993) with discretized signals and no dominance regions.
@@
where $\tau$ is a non-negative integer or infinite ($\tau=\infty$). If $r$ is sufficiently close to one, this game has exactly three pure-strategy BNE: $s^{0}$, $s^{\infty}$, and $s^{\tau^{*}}$, where $\tau^{*}\geq 2$ is the smallest integer satisfying ${r}^{\tau^{*}}<(1-p)/(1+r(1-p))$, provided no integer $\tau$ satisfies this condition with equality. [^13] This condition ensures that type $\tau^{*}$’s best response is action 0 when type $\tau^{*}-1$ chooses action 1 and type $\tau^{*}+1$ chooses action 0. The conditional expected value of the potential function for type $\tau^{*}$ is $r^{\tau^{*}}/(2-p)$ under action 1 and $(1-p)(1-r^{\tau^{*}+1})/(2-p)$ under action 0. If the latter is greater, action 0 is the best response.

### v4

@@ 1 Introduction
Let the analyst’s model be an incomplete information game with two players $1$ and $2$; two actions $\alpha$ and $\beta$ for each player; and two payoff-relevant states $\theta_{1}$ and $\theta_{2}$ occurring with equal probability of 1/21/2.111This game is discussed by Liu (2015). Each player’s set of types is a singleton; that is, it is common knowledge that each player has a uniform prior over $\{\theta_{1},\theta_{2}\}$. The payoffs are summarized in Table 1, where players would be better coordinated by choosing the same actions in state $\theta_{1}$ and the different actions in state $\theta_{2}$. This game has an infinite number of Bayes Nash equilibria (BNE) because the expected payoff for each player is $1/2$ under every action profile.
@@
$\alpha$ 0 $1/4$
$\beta$ $1/4$ 0
Table 2: Joint probabilities over actions and a state
@@
A generalized potential function is associated with belief-invariant correlating devices that assign each player an action subset as a recommendation to choose an action from that subset. An important belief-invariant correlating device is one designed to maximize the expected value of a generalized potential function. For any such correlating device, there exists a BIBCE in which each player chooses an action from the recommended subset. This BIBCE is referred to as a GP-maximizing BIBCE.
@@
As the motivating example illustrates, BNE are vulnerable to belief-invariant correlations in nearby games. This naturally raises the question of under what conditions a BNE is robust. Section 6 shows that if a game is a supermodular potential game with a unique potential-maximizing BNE, then this BNE is robust. As a numerical example, we consider a binary-action supermodular game with two players, which is a global game (Carlsson and van Damme, 1993) with discretized signals and no dominance regions. This game has a robust BNE, where players always choose ex-post risk-dominant actions. Even in the absence of dominance regions, allowing a richer set of perturbations for the incomplete information game leads to the same outcome as the global game selection.
@@
The concept of BIBCE is introduced by Forges (2006) and Lehrer et al. (2010) under the additional restriction that action recommendations cannot depend on the state. A BIBCE is a special case of a BCE, which is introduced by Bergemann and Morris (2013, 2016) as an analyst’s prediction when she cannot rule out players having additional information. The work of Bergemann and Morris (2013, 2016) also supports the use of BIBCE in our paper as a candidate for an analyst’s prediction under insufficient and imprecise information. [^2] For further motivation to study BIBCE and its relationship to BCE, see Liu (2015), Bergemann and Morris (2017), and Bergemann et al. (2024). For its relation to other special cases of BCE, see Bergemann and Morris (2019).
@@ 2.1 Incomplete information games
  • $\Theta=\prod_{i\in I}\Theta_{i}$ is an at most countable set of payoff-relevant states, where player $i$’s payoff function is determined by the $i$-th component $\theta_{i}\in\Theta_{i}$. [^5] The restriction to countable types and payoff-relevant states is necessary for the existence of Bayes Nash equilibria and the comparison with the prior literature. BIBCE exist on uncountable type spaces via an argument of Stinchcombe (2011) as discussed by Bergemann et al. (2024), so our results that restrict attention to BIBCE would extend to uncountable type spaces.
@@
Payoff functions are assumed to be bounded, i.e., $\sup_{i,a,\theta}|u_{i}(a,\theta)|<\infty$. For each $i\in I$, let $T_{i}^{*}\subseteq T_{i}$ and $\Theta_{i}^{*}\subseteq\Theta_{i}$ denote the sets of player $i$’s types and payoff-relevant states on the support of $\pi$, respectively: $T_{i}^{*}=\{t_{i}\in T_{i}\mid\pi(t_{i})>0\}$ and $\Theta_{i}^{*}=\{\theta_{i}\in\Theta_{i}\mid\pi(\theta_{i})>0\}$, where $\pi(t_{i})\equiv\sum_{t_{-i},\theta}\pi(t,\theta)$ and $\pi(\theta_{i})\equiv\sum_{t,\theta_{-i}}\pi(t,\theta)$ are the marginal probabilities. Player $i$’s belief is given by $\pi(t_{-i},\theta|t_{i})\equiv\pi(t,\theta)/\pi(t_{i})$ when his type is $t_{i}\in T_{i}^{*}$.
@@ 2.3 Elaborations
Note that if $(\bar{T},\Theta,\bar{\pi},u)$ is the conjunction of $(T,\Theta,\pi,u)$ and a belief-invariant communication rule $\rho$, it is an elaboration of $(T,\Theta,\pi,u)$ with $\tau_{i}$ and $\mu_{i}$ given by $\tau_{i}(t_{i},m_{i})=t_{i}$ and $\mu_{i}(t_{i},m_{i})=m_{i}$ for all $\bar{t}_{i}=(t_{i},m_{i})\in\bar{T}_{i}$ and $i\in I$. [^8] Equation (1) implicitly requires that $\sum_{\bar{t}\in\tau^{-1}(t)}\rho(\mu(\bar{t})|t,\theta)=1$ for all $(t,\theta)\in T^{*}\times\Theta^{*}$ and $\tau(\bar{T}^{*})=T^{*}$ because $\bar{\pi}(\bar{T}\times\Theta)=1$.
@@
A game $(\bar{T},\Theta,\bar{\pi},u)$ is an elaboration of $(T,\Theta,\pi,u)$ if and only if there exists a mapping $\tau_{i}:\bar{T}_{i}\to T_{i}^{*}$ for each $i\in I$ such that
@@ 2.4 Non-redundant games
A non-redundant game [^10] We can characterize a non-redundant game using belief hierarchies. Let $t_{i}]\equiv\{t_{i}^{\prime}\in T_{i}\mid h_{i}(t_{i})=h_{i}(t_{i}^{\prime})\}$ and $t]=([t_{i}])_{i\in I}$. Consider $(T^{0},\Theta,\pi^{0},u)$ with $T^{0}=\{[t]\mid t\in T\}$ and $\pi^{0}([t],\theta)=\pi(\{(t^{\prime},\theta)\,|\,t^{\prime}\in[t]\})$. Then, it is straightforward to show that $(T^{0},\Theta,\pi^{0},u)$ is a non-redundant game of $(T,\Theta,\pi,u)$. will serve as the analyst’s model in the next section. Liu (2015) shows that if two games have the same set of belief hierarchies and one is non-redundant, then the other is an elaboration of the non-redundant game. [^11] Liu (2015) assumes a finite type space.
@@ 3 Robustness
We further assume that the analyst believes that the true game is in a “neighborhood” of elaborations of the analyst’s model. We call such a nearby game an $\varepsilon$-elaboration, which is approximately an elaboration of $(T,\Theta,\pi,u)$. An elaboration is a special case of an $\varepsilon$-elaboration with $\varepsilon=0$ (see Figure 1).
$\varepsilon$-elaborationselaborationsAnalyst’s model(non-redundant game)
Figure 1: The analyst’s model is nested within elaborations and $\varepsilon$-elaborations.
@@
    for all $\bar{t}_{i}\in\bar{T}_{i}^{\flat}$, where $\tau(\bar{t})=(\tau_{i}(\bar{t}_{i}))_{i\in I}$ and $\tau_{-i}(\bar{t}_{-i})=(\tau_{j}(\bar{t}_{j}))_{j\neq i}$. We call the mapping $\tau$ an elaboration mapping (with some abuse of terminology).
@@
for some $\sigma\in\mathcal{E}$, where $\bar{\sigma}\circ{\bar{\pi}}(a,\bar{t},\theta)=\bar{\sigma}(a|\bar{t},\theta)\bar{\pi}(\bar{t},\theta)$ and $\sigma\circ\pi(a,t,\theta)=\sigma(a|t,\theta)\pi(t,\theta)$. If $\mathcal{E}=\{\sigma\}$ is a singleton, then $\sigma$ is said to be robust.
To interpret (7), recall Lemma 4, which says every 0-elaboration has a BIBCE $\bar{\sigma}$ that is equivalent to a BIBCE $\sigma$ of $(T,\Theta,\pi,u)$, where (4) holds for all $(a,t,\theta)\in A\times T\times\Theta$. Equation (7) means that (4) holds approximately for sufficiently small $\varepsilon>0$ in the case of $\varepsilon$-elaborations.
@@
Given the definition of robustness, the motivating example in Section 1 is summarized as follows and illustrates why we focus on BIBCE in our robustness concept. We have shown that there exists an $\varepsilon$-elaboration with a unique BNE. This BNE is close to the BIBCE in Table 2 when $\varepsilon$ is small, but it is not close to any BNE of the analyst’s model. This implies that the set of BNE for the analyst’s model is not robust, so we adopt BIBCE for the analyst’s model. We also adopt BIBCE for $\varepsilon$-elaborations because there exists an $\varepsilon$-elaboration whose BNE is not the BIBCE in Table 2. For instance, the analyst’s model is a 0-elaboration, where every BNE differs from the BIBCE in Table 2.
@@
In Kajii and Morris (1997), the analyst’s model is a complete information game. A correlated equilibrium is defined to be robust if, for sufficiently small $\varepsilon>0$, every $\varepsilon$-elaboration has a BNE that is close to the correlated equilibrium. See Section 7.1 for further details.
@@ 4.2 Generalized potentials and BIBCE
A generalized potential function (Morris and Ui, 2005) is defined as a function over $\mathcal{A}\times\Theta$, which contains certain information about players’ preferences.
@@
for all $i\in I$, $a\in A$, and $\theta\in\Theta^{*}$. The next lemma shows that a potential function is a special case of a generalized potential function.
@@ 5.1 Robustness of GP-maximizing BIBCE
Every game admits a generalized potential function with the domain $\mathcal{A}=\{\{A\}\}$, in which case $\mathcal{E}^{F}$ is the set of all BIBCE. The theorem implies that if a BIBCE is unique and is the only element of $\mathcal{E}^{F}$, it is robust.
@@ 5.2 Proof of Theorem 1
Now, let $\overline{\mathcal{E}}^{F}$ be the set of all GP-maximizing BIBCE of $(\bar{T},\bar{\Theta},\bar{\pi},\bar{u})$. Then, by Lemma 8, $\{(\bar{T},\bar{\Theta},\bar{\lambda}^{k},\bar{u})\}_{k=1}^{\infty}$ has a sequence of BIBCE $\{\bar{\sigma}^{k}\}_{k=1}^{\infty}$ such that
@@ 6.1 Supermodular potential games
Let $A_{i}$ be linearly ordered with $\geq_{i}$. We write $\geq_{I}$ for the product order: for $a,b\in A$, $a\geq_{I}b$ if and only if $a_{i}\geq_{i}b_{i}$ for all $i\in I$. We say that $(T,\Theta,\pi,u)$ is a supermodular game if, for each $\theta\in\Theta$, the ex-post game with a payoff function profile $(u_{i}(\cdot,\theta))_{i\in I}$ exhibits strategic complementarities; that is, for each $i\in I$ and $a,b\in A$ with $a\geq_{I}b$,
@@
For each belief-invariant decision rule $\sigma\in\Sigma^{BI}$, we define a pure strategy profile $\overline{\sigma}=(\overline{\sigma}_{i})_{i\in I}$, where for each $i\in I$ and $t_{i}\in T_{i}$, player $i$ of type $t_{i}$ chooses the maximum action in the support of $\sigma_{i}(\cdot|t_{i})$; that is, $\bar{\sigma}_{i}(\cdot|t_{i})$ places probability one on $\max\{a_{i}\mid\sigma_{i}(a_{i}|t_{i})=\sigma\left(\{a_{i}\}\times A_{-i}|\left(t_{i},t_{-i}\right),\theta\right)>0\}$. Similarly, we define another pure strategy profile $\underline{\sigma}=(\underline{\sigma}_{i})_{i\in I}$, where player $i$ of type $t_{i}$ chooses the minimum action in the support of $\sigma_{i}(\cdot|t_{i})$.
The following proposition establishes that if $\sigma$ is a P-maximizing BIBCE, then the strategy profiles $\overline{\sigma}$ and $\underline{\sigma}$ are also P-maximizing BIBCE.
@@
As an example of such a BNE, we consider a binary-action supermodular game with two players, each choosing either action 1 or action 0 (which can be interpreted as “invest” and “not invest,” respectively). The set of states is given by $\Theta=\{0,1,2,3,\ldots\}$. A payoff table and a potential function for each state $n\in\Theta$ are given below, where $0<r<1$.
@@
action 1 ${r}^{n+1}$ 0
action 0 0 $1-{r}^{n+1}$
State $n\in\Theta$ occurs with probability $p(1-p)^{n}>0$, where $p\in(0,1)$. Each player has the following information partition of $\Theta$: player $1$’s partition is $\{\{0\},\{1,2\},\{3,4\},\ldots\}$, and player $2$’s partition is $\{\{0,1\},\{2,3\},\{4,5\},\ldots\}$, which are analogous to those in the email game. The type of a player with a partition $\{k-1,k\}$ (or $\{0\}$) is referred to as type $k$ (or type 0). Thus, the sets of player 1’s types and player 2’s types are $T_{1}=\{0,2,4,\ldots\}$ and $T_{2}=\{1,3,5,\ldots\}$, respectively. This game can be interpreted as a global game (Carlsson and van Damme, 1993) with discretized signals and no dominance regions.
@@
where $\tau$ is a non-negative integer or infinite ($\tau=\infty$). If $r$ is sufficiently close to one, this game has exactly three pure-strategy BNE: $^{0}, ss^{\infty}$, and $s^{\tau^{*}}$, where $\tau^{*}\geq 2$ is the smallest integer satisfying ${r}^{\tau^{*}}<(1-p)/(1+r(1-p))$, provided no integer $\tau$ satisfies this condition with equality. [^14] This condition ensures that type $\tau^{*}$’s best response is action 0 when type $\tau^{*}-1$ chooses action 1 and type $\tau^{*}+1$ chooses action 0. The conditional expected value of the potential function for type $\tau^{*}$ is $r^{\tau^{*}}/(2-p)$ under action 1 and $(1-p)(1-r^{\tau^{*}+1})/(2-p)$ under action 0. If the latter is greater, action 0 is the best response.